0.0055 mol of gas undergoes the process 1 rightarr
Answer #1

Using Ideal Gas eq,
p*v = nRT
As you know, p1 = 1 atm
v1 = 100 cm^3
n = 0.0055 mol

p1*v1 = n*RT1
1 * 101325 * 100 * 10^-6 = 0.0055 * 8.314 * T1
T1 = 221.6 K